In the Newsvendor problem, suppose that the purchase quantity is 20 and the demand is 25. The surplus quantity isNewsvendor Problem: How many newspapers to purchase each day? C = cost to purchase a newspaper Q = number of newspapers the v

Answer Options:

0

5

unable to be determined

-5
Answer #1

Surplus Quantity=MAX(0,Q-D)=MAX(0,Number of newspapers purchased-Number of newspapers demanded)=MAX(0,20-25)=0